OpenStudy (anonymous):

Suppose you want to earn $80. What is the difference of the maximum number of hours and minimum number of hours you need to work to make $80 if you are paid at two different rates? rate A = $5/h rate B = $8/hr?

OpenStudy (dtan5457):

Divide each rate by 80.

OpenStudy (dtan5457):

RATE A: 80/5=16 hours RATE B: 80/8=10 hours difference =16-10=6 hour difference

OpenStudy (dtan5457):

@AleCR

OpenStudy (anonymous):

Thanks

OpenStudy (dtan5457):

NP. Feel free to fan and medal for more help :)

OpenStudy (anonymous):

\[\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\]